40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 26 00:15:51 2013, Annalisa, Update, Green Locking, ETMY - green locking and beat note setup 
    Reply  Thu Jun 27 17:41:08 2013, Annalisa, Update, Green Locking, ETMY -beat note found! beat_note.JPG
Message ID: 8768     Entry time: Thu Jun 27 17:41:08 2013     In reply to: 8751
Author: Annalisa 
Type: Update 
Category: Green Locking 
Subject: ETMY -beat note found! 

 

 Y arm beat note found!

Procedure

  • Arm lock on IR to align the mirrors
  • Green Laser locked on the arm
  • Green Transmission on the PSL and PSL green beam aligned into the BeatPD: they have been aligned both in the near field (looking at the beam on the camera) and in the far field (removing the DC PD and looking at the two beams on the wall)
  • Checked the PSL temperature and, following the plot of the beat note measurement between "Alberto" laser and PSL reported in elog 8396, I got an idea of the range of temperature where the beat note could be found (I used the values of the second curve)
  • Scanned the Y-green laser temperature using the slow servo on the ALS command window

Data

  • PSL temperature = 31.58°C
  • PSL slow servo temperature offset = 0
  • "Alberto" laser temperature = 40.35 °C (ADJ = 0)
  • Thermal output offset on the ALS screen = -11140
  • Beat note frequency = 22 MHz
  • Beat note amplitude = -31.7 dBm

The green transmission on the PSL reads about 500 cts, and the transmitted power is about 50 uW.

(the second peak on the screen in the picture is the 29 MHz of the MC)

Attachment 1: beat_note.JPG  6.795 MB  | Hide | Hide all
beat_note.JPG
ELOG V3.1.3-